OpenStudy (anonymous):

Change the given equation into slope-intercept form. 2x + 5y = -10

OpenStudy (sasogeek):

make y the subject of the equation

OpenStudy (anonymous):

so, 5y = 2x + -10

OpenStudy (sasogeek):

2x + 5y = -10 5y = -2x -10 y = -(2/5)x - (10/5) y = -(2/5)x - 2

OpenStudy (anonymous):

thnx